# Need for zkSpeed: Accelerating HyperPlonk for Zero-Knowledge Proofs

### v1

@@ 3.3. HyperPlonk Protocol Steps
3.3.1. Commit Witnesses
The first step in HyperPlonk’s prover involves witness commitments. In this step, we compute MSMs between elliptic curve points and the witness polynomials to reduce each polynomial to individual commitments. The witness polynomials in HyperPlonk are typically ”Sparse”, meaning that 90% of the values are either 0 or 1, and 10% are up to the full scalar width. SZKP (Daftardar et al., 2024) handles this by using Sparse MSMs. We implement Sparse MSMs using the MSM unit for this step.
@@ 3.3.2. Gate Identity
The Gate Identity step confirms that each gate in the circuit performs its computation correctly. The intuition is that if so, then should evaluate to 0 for each gate, equivalently, for each binary assignment of . Further, the sum of should also equal 0, which can be confirmed by running SumCheck on . However, since the prior statement is only a necessary but not sufficient condition for correct computation, HyperPlonk performs SumCheck over a polynomial , where is a multilinear polynomial with coefficients computed from random challenges as shown in Figure 2 (left). This is known as a ZeroCheck. The Gate Identity step relies on three units: the “multifunction tree unit” (Section 4.3) to build , a SumCheck unit, and the MLE Update unit to update MLE tables between SumCheck rounds.
@@ 3.3.3. Wiring Identity
This step verifies that the outputs of each gate are routed correctly to inputs of downstream gates. This is achieved via a PermutationCheck, which is performed by constructing a numerator polynomial that encodes the ”natural order” of the outputs and a denominator polynomial that encodes the permuted ordering, and then checking that , called the fraction polynomial, equals 1. As we will discuss, computing is time consuming, since inverting (or, in effect, performing divisions on) elements of a finite field is difficult. From , a product polynomial is constructed by computing the product of all elements of . The and polynomials are both committed to via MSM computations. A final ZeroCheck is also performed to ensure that the structure imposed upon the and polynomials was not violated during their creation, i.e., the sum of these polynomial should be zero everywhere. All polynomials are stored in the form of MLE tables. The Wiring Identity step uses the construct N&D unit, the FracMLE unit, the Multifunction Tree unit (to construct the MLE and later within the ZeroCheck step), the MSM unit, the SumCheck unit, and the MLE Update unit.
@@ 3.3.4. Batch Evaluations
The SumCheck protocols require evaluating the MLE at a specific point. Batch evaluation reduces the evaluation of multiple polynomials at multiple points into a single polynomial evaluated at a single point. The batch evaluation step consists of querying the input MLEs, as well as and , at a set of 6 points, some of which are derived from round challenges in the SumCheck portion of the Gate and Wiring Identity steps, and some of which are fixed at compile time. In total, 22 total evaluations are performed among 13 polynomials using 6 distinct points. This step is almost the exact reverse of the step in 3.3.2 used to construct , since we are compressing an entire MLE into 1 value, while builds an entire MLE from values. These polynomial queries are then used by the verifier to check that the prover obeyed the protocol. The Batch Evaluation step uses the Multifunction Tree unit to compute each MLE query.
@@ 3.3.5. Polynomial Opening
This step is used to succinctly verify the correctness of the prover’s batch evaluations. For brevity, we will omit details of thepolynomial opening step. To summarize, it involves first computing 6 MLEs as a linear combination of the MLEs from Eq. 1, and . Then, 6 more MLEs are constructed from the query points in Section 3.3.4. These 12 MLEs are combined into a MLE “dot-product”, upon which a final SumCheck is computed. To avoid confusion, we will refer to this final SumCheck as ”OpenCheck” and ”SumCheck” to refer to the underlying computation.
After OpenCheck, the first 6 MLEs used as OpenCheck’s inputs are linearly combined with the OpenCheck’s round challenges to construct a final MLE, which is denoted as . This MLE is first reduced to half its size, and it is used as the scalar set for a -point MSM. We then halve the scalar set and perform MSMs. For example, if , we compute a -point MSM, then a -point MSM, all the way to a -point MSM. The Polynomial Opening uses the MLE combine unit to compute the linear MLE combinations, the Multifunction Tree unit to build the 6 MLEs, the SumCheck unit to perform OpenCheck, and the MSM Unit to compute MSMs.
@@ 3.3.6. SHA3
zkSNARKs are non-interactive and use SHA3 to generate challenges as well as transcripts. (Transcripts are logs of proof-related computations checked by the verifier.) Because SHA3 is used in-between HyperPlonk steps to log computed values, SHA3 acts as an order-enforcing mechanism. This means most of the protocol steps must be executed in series, as shown in Figure 8.
@@ 3.3.7. Compute Demands of HyperPlonk
Table 1 summarizes profiling results to characterize key functions and understand the sources of performance overhead and hardware needs when accelerating HyperPlonk. There are too many functions to list, and we present the twelve with the highest computational density, which is defined as modular multiplications (modmuls) per Byte, as done in prior work (Kim et al., 2022b; de Castro et al., 2021) (note SHA3 has no modmuls). Additionally, the reference CPU implementation is provided as a link. First, we observe that all functions require an immense amount of computation: ranging from millions to billions of 255/351b modmuls (comprising three integer multiplications) over all invocations of each function. For example, the data for Wire Identity MSMs modmuls reflects two function calls; for ZeroCheck Rounds there is one function call (see links in table). This motivates the need for both specialized modmul units and a high degree of parallelism to mitigate overhead. Second, compute intensity drops off sharply after the third function (since data reuse is limited), and the data input/output sizes for all functions are large, typically hundreds of megabytes up to terabytes as problem sizes scale to more gates. This motivates the need for large on-chip scratchpads to mitigate off-chip data movement when possible and high off-chip (i.e., HBM) bandwidth.
@@ 4. HyperPlonk Accelerator Units
HyperPlonk’s protocol is based on the BLS12-381 elliptical curve. Here, all MLE datatypes are 255 bits wide, and all elliptical curve points in the MSMs are 381 bits wide. All MLE and MSM operations involve modular arithmetic primitives. These are built into each accelerator unit that requires them. In total, zkSpeed comprises eight accelerator units, we describe each below.
@@ 4.1. SumCheck and MLE Update
Three HyperPlonk steps use SumCheck: ZeroCheck (Section 3.3.2), PermutationCheck (Section 3.3.3), and OpenCheck (Section 3.3.5). Each step has a unique polynomial shown, respectively, below:
In these equations, is a challenge from the SHA3 unit, and all other symbols represent multilinear polynomials. These polynomials share a common sum-of-products representation, but are each unique and require slightly different datapaths. We develop a unified SumCheck Unit that can handle each of these polynomials. We highlight the key contributions of our architecture next.
@@ 4.1.1. Sumcheck Round PE Microarchitecture
In ZeroCheck and PermutationCheck, there are polynomials that appear multiple times across terms. In HyperPlonk’s CPU baseline, the boolean hypercube summations are performed iteratively term-by-term, incurring redundant computation for these repeating polynomials. We address this by computing all evaluations for each polynomial in parallel. For example, in Equation 2, the polynomial has to be evaluated at and . This computation only needs to be performed once before being used to compute the product in each term of . Figure 3 shows an example of how our Sumcheck PEs handle ZeroCheck computation for . Each unique MLE takes its respective values (e.g., ), and uses these values to construct the needed evaluations. Then, each product in Equation 2 must be computed for each evaluation. The products at the evaluation points are then accumulated into registers. Due to the inherent degree imbalance, some terms have fewer evaluations; the additional evaluations are computed via Barycentric Interpolation (Berrut and Trefethen, 2004). This is omitted in the figure since it only adds a fixed cost at the end of each round (23 modmuls for ZeroCheck and 46 modmuls for PermCheck). Each MLE observes different datapaths, so we use a specialized design to exploit high reuse, full-pipelining, and high levels of parallelism.
@@ 4.1.2. Streaming approach
At the start of the protocol, the MLE tables that are provided to the prover can be stored on-chip. However, as these MLEs undergo rounds of SumCheck, the process of incorporating challenges into the MLE values expands binary values to the full 255 bits. Though the number of MLE table entries reduces by half each round, the data itself grows by over 100, so the total storage cost for storing all MLEs is intractable. However, each round, the intermediate values of MLE tables are only used by the main SumCheck computation and then by the MLE Update to be halved in size. Since there is no data reuse in-between rounds, we adopt a streaming-based solution to alleviate the pressure on on-chip SRAM storage. The key tradeoff here is that our SumCheck and MLE Update units become memory-bound, since each MLE in Equations 2-4 must be updated and written back to off-chip memory after SumCheck rounds, necessitating off-chip traffic. Fortunately, recent advances in high-bandwidth memory (HBM) can supply very high bandwidths to offset this. This is inline with many other cryptographic computing accelerators, which also rely on HBM (Daftardar et al., 2024; Ebel and Reagen, 2024; Samardzic et al., 2021, 2022; Kim et al., 2022b, 2023, a). We analyze the bandwidth sensitivity of the SumCheck computations in our evaluation section.
@@ 4.1.3. Scaling to Multiple PEs
Each SumCheck PE handles the product and sum for one iteration of the boolean hypercube sum. For example, in Figure 1, this corresponds to one red region and the first product in the green region for each . These iterations run in parallel with multiple PEs storing their own accumulation registers. The bottom of Figure 3 shows the accumulation of each evaluation across boolean hypercube values (equivalently, across indices of the MLE tables). After all evaluations are done, an MLE Update PE handles the updates for one MLE, and can provision multiple modmuls. Multiple PEs can run in parallel, handling each MLE independently.
@@ 4.1.4. Unified SumCheck PE
The three polynomials , , have different datapaths. We use HLS to generate a unified PE that handles each SumCheck variation used in HyperPlonk. Each PE requires 94 modular multipliers, compared to 184 modular multipliers without resource sharing, saving 48.9% on area. HyperPlonk’s CPU baseline is designed to support any composition of multilinear polynomials for different protocols, not just the three we have shown. In software, repeating polynomial computations greatly improves programmability as opposed to having specialized functions to handle the specific computation patterns we optimize.
NoCap (Samardzic et al., 2024) is a recent accelerator that also accelerates SumCheck, but there are critical differences at the protocol-level that motivate our architecture. NoCap implements Spartan (Setty, 2020), which uses R1CS encodings resulting in two SumCheck instances that look as follows: and . NoCap uses a vector architecture with 2048 PEs to process boolean hypercube instances with a Beneš network to sum across PEs. This makes sense for NoCap because Spartan’s polynomials are degree and with up to two terms. In contrast, HyperPlonk’s polynomials have a more heterogeneous structure; there are more terms of varying degree. This complexity arises from the usage of the control signal MLEs () to represent gates. These MLEs are required to keep HyperPlonk verifier costs low (in Section 7, we see NoCap’s verifier is slower). Consequently, mapping HyperPlonk’s polynomials to a vector architecture would put more pressure on vector register files because of the high amount of intermediate values needed to be read. As seen in Figure 3, our SumChecks also require very complex communication, which can increase bandwidth pressure and may not be efficient to implement with a Beneš network. Further, our specialized PEs immediately reuse values without relying on register files to store the numerous amount of intermediates. We have fewer, heftier PEs so data movement costs are less relative to those incurred by a vector processor.
@@ 4.2. MSMs
MSMs are used in three Hyperplonk steps: Witness Commitments, the Wiring Identity, and Polynomial Opening. Several prior works accelerated MSMs. We begin with the base MSM design from SZKP (Daftardar et al., 2024), the state-of-the-art for accelerating Groth16, and propose two optimizations for better performance and area efficiency.
In HyperPlonk, the sparse MSMs run in series are on the critical path. (This is unlike prior protocols like Groth16 where their execution could be masked via parallel processing with the dense MSMs.) Consequently, we opt to use the same MSM hardware unit for both sparse and dense MSMs in zkSpeed (as opposed to separate units in SZKP). zkSpeed uses a similar scheme as SZKP for handling sparse computations. First, we compute the sum of all points corresponding to -valued scalars. This is done by fetching the points corresponding to -valued scalars into the MSM unit’s SRAM banks. Then, using a tree-based approach, we feed two points at a time to the pipelined Point Adder (PADD) unit, with the result written back to the SRAM banks. This process repeats until we have reduced all points to the final sum for points corresponding to -valued scalars. Note that in this step, we need not fetch scalars into the MSM’s SRAM banks since they are all . Then, we use Pippenger’s algorithm (Pippenger, 1976) on the remaining 10% of (dense) scalars. For the Dense MSMs in the Wiring Identity and Polynomial Opening steps, we use Pippenger’s algorithm for the full MSM computation.
We improve upon SZKP’s architecture in two ways. First, we note that elliptical curve points, while being three-dimensional, are initialized as coordinates in HyperPlonk. Thus, we only fetch two coordinates per point, saving off-chip bandwidth. We further save on-chip SRAM area. SZKP provisions one scalar memory bank and three point memory banks, to hold coordinates. In Hyperplonk Sparse MSMs do not need to store scalars, however, the tree-based addition partials still need buffering (their coordinates are no longer 1-valued) necessitating a coordinate memory bank. Therefore, zkSpeed allocates three SRAM memory banks. In dense MSM operation, the memory bank is reused to store the scalars, and since partial sums are only stored in bucket registers, we do not need to provision another dedicated memory bank, as in SZKP. This represents a savings of 18% in on-chip SRAM area compared to having a dedicated scalar memory bank.

### v2

@@ 3.3. HyperPlonk Protocol Steps
3.3.1. Witness Commits
The first step in HyperPlonk’s prover involves witness commitments. In this step, we compute MSMs between elliptic curve points and the witness polynomials to reduce each polynomial to individual commitments. The witness polynomials in HyperPlonk are typically “Sparse”, meaning that 90% of the values are either 0 or 1, and 10% are up to the full scalar bitwidth. SZKP (Daftardar et al., 2024) handles this by using Sparse MSMs. We implement Sparse MSMs using the MSM unit for this step.
@@ 3.3.2. Gate Identity
The Gate Identity step confirms that each gate in the circuit performs its computation correctly. The intuition is that if so, then in Equation 1 should evaluate to 0 for each gate, or equivalently, for each binary assignment of . Further, the sum of over all ’s should also equal 0, which can be confirmed by running SumCheck on . However, since the prior statement is only a necessary but not sufficient condition for correct computation, HyperPlonk performs SumCheck over a polynomial , where is a multilinear polynomial with evaluations computed from random challenges (the polynomial of Lemma 2.1 in (Chen et al., 2022)). We will henceforth refer to the construction of such a polynomial as Build MLE. Then, the sum over the boolean hypercube of should also be 0; this is known as a ZeroCheck. The Gate Identity step relies on three units: the Multifunction Tree unit (Section 4.3) to build , the SumCheck unit, and the MLE Update unit to update MLE tables between SumCheck rounds.
@@ 3.3.3. Wiring Identity
This step verifies that the outputs of each gate are routed correctly to inputs of downstream gates. This is achieved via a PermutationCheck (PermCheck), which is performed by constructing a numerator polynomial that encodes the natural ordering of the outputs and a denominator polynomial that encodes the permuted ordering (derived from permutation polynomials ), and then checking that , which we refer to as Fraction MLE, equals 1. As we will discuss, computing is time consuming, since inverting (or, in effect, performing divisions on) elements of a finite field is difficult. The Product MLE, , is generated by computing a cumulative product over to form the first half, then recursively applying cumulative products on itself to complete the second half. Then, the and polynomials are both committed to via MSM computations. A final ZeroCheck is also performed to ensure that the structure imposed upon and was not violated during their creation, i.e., the sum over the boolean hypercube of the resulting constraint polynomial should be zero everywhere. All polynomials are stored in the form of MLE tables. The Wiring Identity step uses the Construct N&D unit, the FracMLE unit, the Multifunction Tree unit (to construct the Product MLE and later for the Build MLE step within ZeroCheck), the MSM unit, the SumCheck unit, and the MLE Update unit.
@@ 3.3.4. Batch Evaluations
The SumCheck protocols require evaluating several MLEs at specific points. This is achieved with batch evaluations, where the input MLEs, as well as and , are queried at 6 points, some of which are derived from round challenges in the SumCheck portion of the Gate and Wiring Identity steps, and some of which are fixed at compile time. In total, 22 total evaluations are performed among 13 polynomials using 6 distinct points. This step is almost the exact reverse of the step in 3.3.2 used to construct , since we are compressing an entire MLE into 1 value, while is an entire MLE built from values. We will henceforth refer to the operation of evaluating an MLE at a point (consisting of field elements) as MLE Evaluate. These polynomial queries are then used by the verifier to check that the prover obeyed the protocol. The Batch Evaluation step uses the Multifunction Tree unit.
@@ 3.3.5. Polynomial Opening
This step succinctly verifies the correctness of the prover’s batch evaluations. For brevity, we will omit details of the polynomial opening step. To summarize, it involves first computing 6 MLEs as a linear combination of the MLEs from Eq. 1 and and . Then, 6 more MLEs are constructed from the query points in Section 3.3.4. These 12 MLEs are linearly combined into a single MLE, upon which a final SumCheck is computed. To avoid confusion, we will refer to this final SumCheck as OpenCheck and use SumCheck to refer to the underlying computation.
After OpenCheck, the first 6 MLEs used as OpenCheck’s inputs are linearly combined with the OpenCheck’s round challenges to construct a final MLE, which is denoted as . This MLE is first reduced to half its size, and it is used as the scalar set for a -point MSM. We then halve the scalar set and perform MSMs. For example, if , we compute a -point MSM, then a -point MSM, all the way to a -point MSM. The Polynomial Opening step uses the MLE Combine unit to compute the linear MLE combinations, the Multifunction Tree unit to build the 6 MLEs, the SumCheck unit to perform OpenCheck, and the MSM Unit to compute MSMs.
@@ 3.3.6. SHA3
zkSNARKs are non-interactive and use SHA3 to generate challenges as well as maintain transcripts. Transcripts are logs of proof-related computations checked by the verifier. In HyperPlonk, SHA3 is invoked between steps to “log” computed values, which is achieved by updating the SHA3 state. Intuitively, this ensures that all future challenges are bound to the existing history of computations logged in the transcript. As a result, SHA3 effectively acts as an order-enforcing mechanism. This means the protocol steps must be executed in series, as shown in Figure 2.
@@ 3.3.7. Compute Demands of HyperPlonk
Table 1 summarizes profiling results to characterize key functions and understand the sources of performance overhead and hardware needs when accelerating HyperPlonk. There are too many functions to list, so we present the twelve with the highest arithmetic intensity, which is defined as modular multiplications (modmuls) per byte, as done in prior work (Kim et al., 2022b; de Castro et al., 2021) (note SHA3 has no modmuls). Additionally, the reference CPU implementation is provided as a link. First, we observe that all functions require an immense amount of computation: ranging from millions to billions of 255/381-bit modmuls (comprising three integer multiplications) over all invocations of each function. For example, the data for Wire Identity MSMs modmuls reflects two function calls; for ZeroCheck Rounds there is one function call (see links in table). This motivates the need for both specialized modmul units and a high degree of parallelism to mitigate overhead. Second, compute intensity drops off sharply after the third function (since data reuse is limited), and the data input/output sizes for all functions are large, typically hundreds of megabytes up to terabytes as problem sizes scale to more gates. This motivates the need for large on-chip scratchpads to mitigate off-chip data movement when possible and high off-chip (i.e., HBM) bandwidth.
@@ 4. HyperPlonk Accelerator Units
HyperPlonk’s protocol is based on the BLS12-381 elliptical curve. Here, all MLE datatypes are 255 bits wide and all elliptical curve points in the MSMs are 381 bits wide. All MLE and MSM operations involve modular arithmetic primitives. These are built into each accelerator unit that requires them. zkSpeed comprises eight accelerator units and we describe each below.
@@ 4.1. SumCheck and MLE Update
Three HyperPlonk steps use SumCheck: ZeroCheck (Section 3.3.2), PermCheck (Section 3.3.3), and OpenCheck (Section 3.3.5). Each step has a unique polynomial shown, respectively, below:
In these equations, is a challenge from the SHA3 unit, and all other symbols represent multilinear polynomials. and are all Built MLEs constructed from SHA3 challenges (see the discussion of in Section 3.3.2). These SumCheck polynomials share a common sum-of-products representation, but are each unique and require slightly different datapaths. We develop a unified SumCheck unit that can handle each of these polynomials and highlight the key contributions of our architecture next.
@@ 4.1.1. SumCheck Round PE Microarchitecture
In ZeroCheck and PermCheck, there are polynomials that appear multiple times across terms. In HyperPlonk’s CPU baseline, the boolean hypercube summations are performed iteratively term-by-term, incurring redundant computation for these repeating polynomials. We address this by computing all evaluations for each polynomial in parallel. For example, in Equation 3, the polynomial has to be evaluated at and because the largest-degree term it appears in has a degree of four. This computation only needs to be performed once before being used to compute the product in each term of . Figure 4 shows an example of how our SumCheck PEs handle ZeroCheck computation for . Each unique polynomial takes its respective values (e.g., in Figure 4), and uses these values to extend to the needed evaluations. Then, each product in Equation 3 must be computed for each evaluation. The products at the evaluation points are then accumulated into registers. Due to the inherent degree imbalance, some terms have fewer evaluations; the additional evaluations are computed via Barycentric Interpolation (Berrut and Trefethen, 2004). This is omitted in the figure since it only adds a fixed cost at the end of each round (23 modmuls for ZeroCheck and 46 modmuls for PermCheck). Each polynomial observes a different datapath, so we use a specialized design to exploit high reuse, full-pipelining, and high levels of parallelism.
@@ 4.1.2. Streaming approach
At the start of the protocol, the MLE tables corresponding to polynomials in Equation 1 are all provided to the prover and can be stored on-chip. However, as these MLEs undergo rounds of SumCheck, the process of incorporating challenges into the MLE tables expands binary values to full 255-bit values. Though the number of MLE table entries reduces by half each round, the data itself grows by over 100 between rounds 1 and 2, so the total storage cost for storing all MLEs is intractable. However, each round, the intermediate values of MLE tables are only used by the main SumCheck computation and then by the MLE Update unit to be halved in size. Since there is no data reuse in-between rounds, we adopt a streaming-based solution to alleviate the pressure on on-chip SRAM storage. The key tradeoff here is that our SumCheck and MLE Update units become memory-bound, since each MLE in Equations 3-5 must be updated and written back to off-chip memory after SumCheck rounds, necessitating off-chip traffic. Fortunately, recent advances in high-bandwidth memory (HBM) can supply very high bandwidths to offset this. This is inline with many other cryptographic computing accelerators, which also rely on HBM (Daftardar et al., 2024; Ebel and Reagen, 2024; Samardzic et al., 2021, 2022; Kim et al., 2022b, 2023, a). We analyze the bandwidth sensitivity of the SumCheck computations in our evaluation section.
@@ 4.1.3. Scaling to Multiple PEs
Each SumCheck PE handles the product and sum for one iteration over the boolean hypercube. For example, in Figure 1, this corresponds to one red region per polynomial (MLE table) and the first product in the green region for each . These iterations run in parallel with multiple PEs storing their own accumulation registers. The bottom of Figure 4 shows the accumulation of each evaluation across boolean hypercube values (equivalently, across indices of the MLE tables). After all evaluations are done, an MLE Update PE handles the updates for one MLE table and can provision multiple modmuls. Multiple PEs can run in parallel, handling MLE tables independently.
@@ 4.1.4. Unified SumCheck PE
The three polynomials , , have different datapaths. We use HLS tools to generate a unified PE that handles each SumCheck variation used in HyperPlonk. Each PE requires 94 modular multipliers, compared to 184 modular multipliers without resource sharing, saving 48.9% on area.
4.1.5. Comparison to Prior SumCheck Implementations
HyperPlonk’s CPU library is designed to support any composition of multilinear polynomials for different protocols, not just the three we have shown. Consequently, repeating polynomial computations as observed in the CPU implementation greatly improves programmability, as opposed to having specialized functions to handle the specific computation patterns we optimize for.
NoCap (Samardzic et al., 2024) is a recent accelerator that also accelerates SumCheck, but there are critical differences at the protocol-level that motivate our architecture. NoCap implements Spartan (Setty, 2020), which uses R1CS encodings resulting in two SumCheck instances that look as follows: and . NoCap uses a vector architecture with 2048 PEs to process boolean hypercube instances with a Beneš network to sum across PEs. This makes sense for NoCap because Spartan’s polynomials are degree and with up to two terms. In contrast, HyperPlonk’s polynomials have a more heterogeneous structure; there are more terms of varying degree. This complexity arises from the usage of the control and constant polynomials () to represent gates. These are required to keep HyperPlonk verifier costs low (in Section 7, we see NoCap’s verifier is slower). Consequently, mapping HyperPlonk’s polynomials to a vector architecture would put more pressure on vector register files because of the high amount of intermediate values needed to be read. As seen in Figure 4, our SumChecks also require very complex communication, which can increase bandwidth pressure and may not be efficient to implement with a Beneš network. Further, our specialized PEs immediately reuse values without relying on register files to store the numerous amount of intermediates. We have fewer, heftier PEs, so data movement costs are less relative to those incurred by a vector processor.
@@ 4.2. MSMs
MSMs are used in three HyperPlonk steps: Witness Commits, Wiring Identity, and Polynomial Opening. Several prior works accelerated MSMs; we begin with the base MSM design from SZKP (Daftardar et al., 2024), the state-of-the-art for accelerating Groth16, and propose two optimizations for better performance and area efficiency.
In HyperPlonk, the Sparse MSMs run in series and are on the critical path. This is unlike prior protocols like Groth16 where their execution could be masked via parallel processing with the Dense MSMs. Consequently, we opt to use the same MSM hardware unit for both Sparse and Dense MSMs in zkSpeed (as opposed to separate units in SZKP). zkSpeed uses a similar scheme as SZKP for handling sparse computations. First, we compute the sum of all points corresponding to -valued scalars. This is done by fetching the points corresponding to -valued scalars into the MSM unit’s SRAM banks. Then, using a tree-based approach, we feed two points at a time to the pipelined Point Adder (PADD) unit, with the result written back to the SRAM banks. This process repeats until we have reduced all points to the final sum for points corresponding to -valued scalars. Note that in this step, we need not fetch scalars into the MSM’s SRAM banks since they are all . Then, we use Pippenger’s algorithm (Pippenger, 1976) on the remaining 10% of (dense) scalars. For the Dense MSMs in the Wiring Identity and Polynomial Opening steps, we use Pippenger’s algorithm for the full MSM computation. We now discuss the two improvements to SZKP’s MSM architecture.
4.2.1. Reduced Memory Footprint
First, we note that elliptical curve points, while being three-dimensional, are initialized as , 1) coordinates in HyperPlonk. Thus, we only fetch two coordinates per point, saving off-chip bandwidth. We further save on-chip SRAM area. SZKP provisions one scalar memory bank and three point memory banks, to hold coordinates. While Sparse MSMs need not store the 1-valued scalars, the tree-based addition partials still need buffering (their coordinates are no longer 1-valued) necessitating the coordinate memory bank. Therefore, zkSpeed allocates three SRAM memory banks. In Dense MSM operation, the memory bank is reused to store the scalars, and since partial sums are only stored in bucket registers, we do not need to provision another dedicated memory bank, as in SZKP. This represents a savings of 18% in on-chip SRAM area compared to having a dedicated scalar memory bank.
4.2.2. Faster Bucket Aggregation